### v1

# Life at low Reynolds number isn’t such a drag

Abstract
The following is an unedited version of two short articles that are forthcoming in Nature Chemical Engineering. Inspired by Purcell’s classic lecture “Life at low Reynolds number”, I discuss how scaling arguments, dimensional analysis, and fundamental concepts from chemical engineering science can be used to quantitatively describe microbial swimming—thereby helping to better understand biological systems and inspiring new engineering advances in turn.
(I) Individuals. Imagine swimming in a pool on a hot summer’s day. What distance $\delta_{\mathrm{macro}}$ do you coast after each stroke that propels you at a speed $V$? The drag force $f_{\mathrm{drag}}\sim\rho V^{2}L^{2}$ exerted by the surrounding fluid causes you to decelerate by an amount $\sim V^{2}/\delta_{\mathrm{macro}}$; here, $L$ represents your overall size and $\rho\sim m/L^{3}$ is your mass ($m$) density. Thus, Newton’s second law dictates that $f_{\mathrm{drag}}\sim mV^{2}/\delta_{\mathrm{macro}}$, or, $\delta_{\mathrm{macro}}\sim L$. That is, you coast a distance comparable to your size, on the order of a meter.
How about for swimming microbes, which are much smaller and slower than us? In this case, fluid dynamics tells us that $f_{\mathrm{drag}}\sim\eta VL$, where $\eta$ is the fluid dynamic viscosity. Newton’s second law then says that $\eta VL\sim mV^{2}/\delta_{\mathrm{micro}}$, or, $\delta_{\mathrm{micro}}\ \sim\ mV/(\eta L)$. This coasting distance can be much smaller than the size of the microbe itself: for the commonly-studied bacterium E. coli in water, $m\simeq 1$ pg, $V\simeq 10$ $\mu$m/s, $\eta\simeq 1$ mPa-s, $L\simeq 1$ $\mu$m, yielding a coasting distance of 0.1 Angstroms — a tenth the size of a hydrogen atom! The viscous drag of the surrounding fluid plays such a dominant role for microbes that it stops them from coasting essentially instantaneously. An interesting consequence of this fact is that microbes cannot use reciprocal motion—i.e., motion that looks the same when played in reverse, as in the case of a scallop repeatedly opening and closing its shell [^1] The scallop is just an illustrative example; in reality, it likely closes its shell sufficiently fast so that $\mathrm{Re}>1$.—to translate large distances. They would just keep going back and forth and wouldn’t get anywhere! Researchers are working to engineer systems that break this so-called “scallop theorem”, [^2] E. Lauga, “Life around the scallop theorem”, Soft Matter, 7, 3060 (2011). potentially helping to establish new classes of self-propelling micromachines.
Taking the ratio between the macroscopic and microscopic coasting distances highlights this increasing relative importance of viscous drag at small length scales, yielding a dimensionless parameter that chemical engineers know well: the Reynolds number ${\mathrm{Re}=\delta}_{\mathrm{micro}}/\delta_{\mathrm{macro}}=mV/\left({\eta L}^{2}\right)=\rho VL/\eta$. As illustrated by the simple calculations above, life at low Reynolds number ($\mathrm{Re}<1$) is very different from the world of large Reynolds number ($\mathrm{Re}>1$) that we humans typically inhabit.
These calculations were beautifully described by Edward Purcell, winner of the Nobel Prize in Physics for discovering nuclear magnetic resonance, in a 1976 lecture to a distinguished group of physicists titled “Life at low Reynolds number”. [^3] E. M. Purcell, “Life at low Reynolds number”, Am. J. Phys., 45, 3 (1977). While his lecture was important in establishing an intellectual framework for thinking about the physics of swimming microbes, it could have arguably been a chemical engineering lecture instead: after all, it uses scaling arguments and ideas from dimensional analysis to compare key length, time, and force scales—just as good chemical engineers do. Indeed, as I highlight below, the contemporary research areas that have emanated from Purcell’s original lecture apply fundamental concepts from chemical engineering science—mass and energy balances, thermodynamics, transport phenomena [^4] J. Happel and H. Brenner, Low Reynolds number hydrodynamics. Prentice Hall, Englewood Cliffs, NJ (1967)., and chemical kinetics—to shed new light on problems at the interface of engineering, biology, chemistry, and physics, and in turn, seek inspiration from these fields to engineer new forms of synthetic microswimmers for e.g., drug delivery and environmental sensing. They are therefore of direct interest to modern-day researchers in chemical engineering and represent an important frontier for the future of the field.
For example: What if the microbe in our thought experiment was swimming in a polymeric fluid instead of just water? This is a question of clear biological importance: microbes often inhabit, and cause infections in, polymeric fluids such as mucus in the lungs and cervicovaginal tract. In this case, the swimming microbe generates an elastic stress $\sim\ \eta\lambda(V/L)^{2}$ in addition to the viscous stress $\sim\ \eta V/L$; here, $\lambda$ is the time scale over which elastic stresses relax in such viscoelastic fluids and $V/L$ characterizes the fluid strain rate generated by swimming. Comparing these stresses yields a new dimensionless parameter: the Weissenberg number $\mathrm{Wi}=\lambda V/L$ [^5] R. J. Poole, “The Deborah and Weissenberg numbers”, The British Society of Rheology, Rheology Bulletin, 53, 32 (2012). When $\mathrm{Wi}>1$, swimming is fast enough to stretch the flexible polymers in the fluid. The resistance generated by polymer elasticity feeds back on the microbe and can modify how it swims in unexpected ways—for example, causing it to wobble less and speed up. [^6] A. Martínez-Calvo, C. Trenado-Yuste, S. S. Datta, “Active transport in complex environments”, in Out-of-Equilibrium Soft Matter: Active Fluids, RSC press (2023). Continuing to explore the coupling between fluid rheology and micro-scale swimming, both for biological microbes as well as engineered microswimmers, is a vibrant area of current research at the interface of fluid dynamics, biology, and the physics of actively moving systems (“active matter”).
One may ask: Why do microbes swim in the first place? The answer may seem obvious: It helps them take up food molecules from their surroundings, like little Pac-Men! But ideas from chemical engineering can help examine the validity of this assumption more quantitatively, as Berg & Purcell did in a follow-up 1977 paper. [^7] H. C. Berg and E. M. Purcell, “Physics of chemoreception”, Biophysical journal, 20, 193, (1977). As a microbe swims through fluid, it drags the fluid within a distance $\sim d$ from its surface along with it — reflecting the no-slip condition of fluids at solid surfaces. So, surrounding nutrient molecules get pushed aside as well, and can only reach the surface of the cell, where they can be taken up, via thermal diffusion with a diffusivity $D$. This process takes a considerable amount of time ${\sim d}^{2}/D$ compared to that of simple advective transport, $\sim d/V$. Comparing these time scales yields a new dimensionless parameter: the Péclet number $\mathrm{Pe}=dV/D$. Berg & Purcell’s numerical calculations of the flow field around an idealized spherical microbe, which built on earlier calculations by chemical engineers, [^8] S. K. Friedlander, “Mass and heat transfer to single spheres and cylinders at low Reynolds numbers”, AIChE journal 3, 43 (1957)., [^9] A. Acrivos and T. D. Taylor, “Heat and mass transfer from single spheres in Stokes flow”, Physics of Fluids 5, 387 (1962). showed that the thickness $d$ of the region surrounding the cell that has $\mathrm{Pe}<1$, through which nutrient transport is diffusion-limited, does not shrink much as swimming speed increases. Thus, they concluded that “in a uniform medium motility cannot significantly increase the cell’s acquisition of material” — suggesting that microbial swimming has a more complex biological function. [^10] A caveat is that, when nutrients are non-uniformly distributed, cells can use chemotaxis to swim up gradients to find more nutrients. Ongoing research is building on these results to consider more complex microbial shapes and swimming behaviors, not just in the context of nutrient acquisition, but also interactions with other suspended materials such as antimicrobials. [^11] C. Lohrmann, C. Holm, S. S. Datta, “Influence of bacterial swimming and hydrodynamics on infection by phages”, Soft Matter 20, 4785 (2024).
(II) Collectives. Our discussion thus far has focused on individual microswimmers. However, in nature and many engineering applications, microbes (and other microswimmers) move in collectives in which they alter each other’s motion via the fluid they push aside, as well as by simply bumping into each other. What are the emergent consequences of these interactions, and how can they be quantitatively described? There is tremendous current interest in addressing this question, [^12] M. C. Marchetti, J. F. Joanny, S. Ramaswamy, T. B. Liverpool, J. Prost, M. Rao, R. A. Simha, “Hydrodynamics of soft active matter”, Reviews of modern physics, 85, 1143 (2013). although its roots date back to a classic 1969 paper by the renowned chemical engineers Bruce Finlayson and L. E. “Skip” Scriven in which they explored the motion of fluids due to “active stresses” generated by, or in, living organisms such as slime molds. [^13] B. A. Finlayson, and L. E. Scriven, “Convective instability by active stress”, Proceedings of the Royal Society of London. A. Mathematical and Physical Sciences 310, 183 (1969).
The microbe from our thought experiment, for example, generates a force equal and opposite to $f_{\mathrm{drag}}\sim\eta VL$ to swim (so-called “force-free” motion at low Reynolds number). Say it swims for a duration $\tau_{R}$ (typically $\sim 10$ s) before changing direction either by choice or inevitable Brownian reorientation. The corresponding amount of work done is then $\sim\eta V^{2}L\tau_{R}$ —approximately 1000 times larger than thermal energy, $k_{B}T$, at room temperature! Hence, for a collective of microswimmers at a dilute number density $n$, by analogy to the Brownian osmotic pressure $\Pi\sim nk_{B}T$, one can define an additional “swim pressure” $\Pi^{\mathrm{swim}}\sim n\cdot\eta V^{2}L\tau_{R}$. (It is tempting to define an “effective temperature” $T_{\mathrm{eff}}\ \sim\eta V^{2}L\tau_{R}/k_{B}$ for this system, but it remains unclear if and when such a description is appropriate in actively moving collectives.) Just as in the kinetic theory of gases, molecular collisions on surrounding boundaries generate our familiar notion of pressure, the collisions between microswimmers and boundaries generates this new kind of active pressure. [^14] S. C. Takatori and J. F. Brady, “Forces, stresses and the (thermo?) dynamics of active matter”, Current Opinion in Colloid & Interface Science, 21, 24 (2016).
Fascinating new effects can emerge in even denser collectives of active microswimmers. For example, when swimming is highly persistent — i.e., the duration $\tau_{R}$ is long compared to the time needed to swim a body length, $L/V$, or equivalently, the reorientation Péclet number ${\mathrm{Pe}}_{R}=L/(V\tau_{R})$ is small — the microswimmers impede each other’s motion, causing a collective to undergo “motility-induced phase separation” (MIPS) into dense and dilute phases. We learn in undergraduate thermodynamics that phase separation in “passive” equilibrium systems typically requires attractive interactions between the constituents; in stark contrast, this active form of phase separation arises from purely repulsive interactions. And remarkably, despite being out of thermodynamic equilibrium, MIPS can be mathematically described using models inspired by the phase separation of thermally equilibrated passive constituents. [^15] M. E. Cates & J. Tailleur, “Motility-induced phase separation”, Annu. Rev. Condens. Matter Phys., 6, 219 (2015). Further exploring how ideas from equilibrium thermodynamics and statistical mechanics can—or cannot—be used to describe active collectives, as well as using these insights to develop ways to e.g., control their spatial organization and put them to work for new forms of actuation, continues to be an exciting frontier of research.
We also learn as undergraduates that fluid flows at low Reynolds number are laminar. Despite living at low Reynolds number, dense collectives of actively moving microswimmers can overturn this paradigm. A prominent example is the phenomenon of “active turbulence”. Because many microswimmers (like bacteria, sperm, and even motor-driven biopolymer assemblies) are elongated in shape, they tend to align when close enough to each other—just like the molecules in a nematic liquid crystal. [^16] D. L. Koch & G. Subramanian, “Collective hydrodynamics of swimming microorganisms: living fluids”, Annual Review of Fluid Mechanics, 43, 637 (2011). Deviations from this alignment between neighbors are penalized, as described by a Frank elastic constant $K$. As a result, microswimmers in a dense collective often move in coherent swirls that, as they interact, give rise to large-scale chaotic flows reminiscent of turbulence! Ongoing research continues to explore the similarities and differences between this intriguing phenomenon and typical inertial turbulence. For example, just as the transition from laminar to inertial turbulence can be parameterized using the Reynolds number, recent work hints that a distinct dimensionless parameter — the “activity number” $A$ — may similarly described the transition to active turbulence. [^17] R. Alert, J. Casademunt, J. F. Joanny, “Active turbulence”, Annual Review of Condensed Matter Physics, 13, 143 (2022). One way to define this parameter is as $A=L_{s}/\sqrt{K/\zeta}$, which compares the system size $L_{s}$ to the length scale $\sqrt{K/\zeta}$above which active swimming overcomes the restoring force of nematic elasticity; here, $\zeta$ is an active stress coefficient, which can be thought of as being $\sim\Pi^{\mathrm{swim}}$, for example. In addition to unraveling these fundamental puzzles, an important direction for chemical engineering research is to develop ways to control active turbulence and harness it for e.g., improved solute mixing.
In closing, I note that our discussion thus far has focused on the fascinating behaviors of microbes swimming in bulk fluids at low Reynolds number. But in many cases, microbes instead swim through elastic solids, such as the extracellular polymer matrix in a bacterial biofilm. Not only does confinement by the elastic matrix change how microbes swim, [^18] T. Bhattacharjee & S. S. Datta, “Bacterial hopping and trapping in porous media”, Nature Communications 10, 2075 (2019)., [^19] S. E. Spagnolie and P. T. Underhill, “Swimming in complex fluids”, Annual Review of Condensed Matter Physics, 14, 381 (2023). but the active stress generated by swimming can deform the matrix itself—unexpectedly leading to formation of self-sustained elastic standing waves. [^20] H. Xu, Y. Huang, R. Zhang, Y. Wu, “Autonomous waves and global motion modes in living active solids”, Nature Physics, 19, 46 (2023). One might imagine that the dimensionless parameter $\Pi^{\mathrm{swim}}/G$, where $G$ is the shear modulus of the elastic matrix, characterizes the onset of this behavior, although such a parameter has not, to my knowledge, yet been identified. Investigating the unusual properties of such “active solids” continues to be a fascinating direction for ongoing research.
The eminent chemical engineer Rutherford “Gus” Aris wrote in 1973: [^21] R. Aris, “Some Interactions between Problems in Chemical Engineering and the Biological Sciences”, in Environmental Engineering: A Chemical Engineering Discipline, pp. 215-225, Dordrecht: Springer Netherlands (1973). “…let us get acquainted with our colleagues working in physiology, biology, zoology, genetics, microbiology and like disciplines for though their problems are infinitely more complex than engineering ones they offer a challenging and important field for our efforts.” Over a half century later, these words ring truer than ever. As the discussion above hopefully highlights, the tools and perspective of chemical engineering have immense potential to help shed new light on biological systems, such as swimming microbes; and these systems in turn present opportunities for new fundamental discoveries and technological advances in chemical engineering. Life at low Reynolds number, it turns out, isn’t such a drag.

### v2

@@
 One way to define this parameter is as $A=L_{s}/\sqrt{K/\zeta}$, which compares the system size $L_{s}$ to the length scale $\sqrt{K/\zeta}$above which active swimming overcomes the restoring force of nematic elasticity; here, $\zeta$ is an active stress coefficient, which can be thought of as being $\sim\Pi^{\mathrm{swim}}$, for example.
+[^18] A related parameter is the Ericksen number, $\mathrm{Er}=\eta VL/K$, which compares the force generated by swimming $\eta VL$ to the elastic restoring force $K$.
 In addition to unraveling these fundamental puzzles, an important direction for chemical engineering research is to develop ways to control active turbulence and harness it for e.g., improved solute mixing.
@@
 But in many cases, microbes instead swim through elastic solids, such as the extracellular polymer matrix in a bacterial biofilm.
-Not only does confinement by the elastic matrix change how microbes swim, [^18] T.
+In this case, confinement by the elastic matrix changes how microbes swim.
+[^19] T.
 Bhattacharjee & S.
 S.
-Datta, “Bacterial hopping and trapping in porous media”, Nature Communications 10, 2075 (2019)., [^19] S.
+Datta, “Bacterial hopping and trapping in porous media”, Nature Communications 10, 2075 (2019).
+This change can be described by two new dimensionless parameters: the Deborah number $\mathrm{De}=t_{\mathrm{mat}}/t_{\mathrm{swim}}$ comparing the relaxation time scale of the matrix $t_{\mathrm{mat}}$ to a characteristic time scale of swimming $t_{\mathrm{swim}}$, and the Benes number $\mathrm{Bn}=L_{\mathrm{mat}}/L$ comparing the confinement length scale imposed by the matrix $L_{\mathrm{mat}}$ to the size of the microbe $L$ [^20] S.
 E.
@@
 Underhill, “Swimming in complex fluids”, Annual Review of Condensed Matter Physics, 14, 381 (2023).
-but the active stress generated by swimming can deform the matrix itself—unexpectedly leading to formation of self-sustained elastic standing waves.
+But also, the active stress generated by swimming can deform the matrix itself—unexpectedly leading to formation of self-sustained elastic standing waves.
 [^21] H.